# Distribution of mixed character sums and extremal problems for Littlewood polynomials

### v1

@@ 8.2. Proof of Corollary 2
For $\alpha=.2$ and $\beta=1.1$, for all large enough primes $q$ and all $k\in\{0,1,\ldots,q-1\}$ we have
@@
The proof is identical to the proof of Lemma 10, provided we can show the additional term $\pm e(-.2k+t)/q^{1/2}$ does not affect the lower bounds in our differential inequalities significantly. In fact, we will show that for the same five intervals covering $[0,1]$ as in the previous subsection, in each interval either the real part or the imaginary part of $H^{\pm}_{k,p}$ satisfies exactly the same linear differential inequality. It is clear that it suffices to show that the first three derivatives (in $t$) of $y(t)=e(-.2k+t)/q^{1/2}$ are small. Indeed, we uniformly have $y^{(l)}(t)=O_{l}(1/q^{1/2})$ for all $l\in\mathbb{N}$. We then apply Proposition 8 and finish in exactly the same way as before. ∎
Note that for the continuous functional $\ell_{\varepsilon}$, the additional term $e(-.2k+t)/q^{1/2}$ converges to 0 uniformly in $t$, therefore we have
$\lim_{q\to\infty}\mathbb{E}\int_{0}^{1}\log|{H}^{\pm}_{q}(t)|\mathds{1}_{\lvert{H}^{\pm}_{q}(t)\rvert\geq\varepsilon}\text{d}t=\mathbb{E}\int_{0}^{1}\log|{G_{.2,1.1}}(t)|\mathds{1}_{|{G_{.2,1.1}}(t)|\geq\varepsilon}\text{d}t$
by Theorem 2. We conclude in exactly the same way as in the previous subsection, i.e. we have
(10) $\lim_{q\to\infty}\mathbb{E}\int_{0}^{1}\log\left|\frac{1}{1.1^{1/2}}{H}^{\pm}_{q}(t)\right|\text{d}t=\mathbb{E}\int_{0}^{1}\log\left|\frac{1}{1.1^{1/2}}{G_{.2,1.1}}(t)\right|\text{d}t.$
To finish the proof of Corollary 2, note that the left hand-side of (10) is the limit of the logarithm of normalized Mahler measures of the Littlewood polynomials $\sum_{n=1}^{\lfloor 1.1q\rfloor}\chi(n+\lfloor.2q\rfloor)x^{n}\pm x^{q-\lfloor.2q\rfloor}$ (upon noticing the Mahler measure is invariant under multiplication by powers of $x$).
@@ 9. Proof of Theorem 4
$F_{p,a}(t)=\sum_{n\leq p}\left(\frac{n+a}{p}\right)e(nt).$
@@
$\lim_{p\to\infty}\frac{1}{\sqrt{p}}\|F_{a,p}(t)\|_{2k}=\phi_{k}(\alpha)$
when $a/p\to\alpha$. We will show that $\phi_{k}(\alpha)$ achieves its minimum at $\alpha=1/4$ for every $k\in\mathbb{N}$.
We now begin the proof. By Theorem 2, we have
$\phi_{k}(\alpha)=\frac{1}{(2\pi)^{2k}}\int_{0}^{1}\mathbb{E}\left\lvert(e(t)-1)\sum_{m\in\mathbb{Z}}\frac{e(m\alpha)}{m+t}\mathbb{Y}(m)\right\rvert^{2k}\text{d}t$
where the $\mathbb{Y}(m)$ are $\pm 1$-valued uniform random variables. Expanding the sum, the integral is equal to
$\int_{0}^{1}\lvert e(t)-1\rvert^{2k}\sum_{m_{i},n_{i}\in\mathbb{Z}}\Bigg(\frac{e(\alpha(m_{1}+\ldots+m_{k}-n_{1}-\ldots-n_{k}))}{(m_{1}+t)\cdots(m_{k}+t)(n_{1}+t)\cdots(n_{k}+t)}\times\\ \mathbb{E}\big[\mathbb{Y}(m_{1})\cdots\mathbb{Y}(m_{k})\mathbb{Y}(n_{1})\cdots\mathbb{Y}(n_{k})\big]\Bigg)\text{d}t.$
@@
$\int_{0}^{1}\mathbb{E}\left\lvert(e(t)-1)\sum_{m\in\mathbb{Z}}\frac{e(m\alpha)}{m+t}\mathbb{Y}(m)\right\rvert^{2k}\text{d}t=\\ \int_{0}^{1}\lvert e(t)-1\rvert^{2k}\sum_{r=0}^{k}\sum_{\begin{subarray}{c}m_{2r+1},\ldots,m_{k}\in\mathbb{Z}\\ n_{2r+1},\ldots,n_{k}\in\mathbb{Z}\end{subarray}}\frac{\mathbb{E}\big[\mathbb{Y}(m_{2r+1})\cdots\mathbb{Y}(m_{k})\mathbb{Y}(n_{2r+1})\cdots\mathbb{Y}(n_{k})\big]}{(m_{2r+1}+t)\cdots(m_{k}+t)(n_{2r+1}+t)\cdots(n_{k}+t)}M_{2r,t}(\alpha)\text{d}t,$
@@
Now writing $\theta=2\alpha$ and continuing from Equation (11), we write the square of the absolute value as the sum of the squares of the real and imaginary parts to get
@@ Acknowledgements
We are grateful to Yu Chen Sun and Victor Wang for fruitful conversations. B.Sh. is funded by a University of Bristol PhD scholarship.

### v2

@@ 8.2. Proof of Corollary 2
For $\alpha=0.2$ and $\beta=1.1$, for all large enough primes $q$ and all $k\in\{0,1,\ldots,q-1\}$ we have
@@
The proof is identical to the proof of Lemma 10, provided we can show the additional term $\pm e(-0.2k+t)/q^{1/2}$ does not affect the lower bounds in our differential inequalities significantly. In fact, we will show that for the same five intervals covering $[0,1]$ as in the previous subsection, in each interval either the real part or the imaginary part of $H^{\pm}_{k,p}$ satisfies exactly the same linear differential inequality. It is clear that it suffices to show that the first three derivatives (in $t$) of $y(t)=e(-0.2k+t)/q^{1/2}$ are small. Indeed, we uniformly have $y^{(l)}(t)=O_{l}(1/q^{1/2})$ for all $l\in\mathbb{N}$. We then apply Proposition 8 and finish in exactly the same way as before. ∎
Note that for the continuous functional $\ell_{\varepsilon}$, the additional term $e(-0.2k+t)/q^{1/2}$ converges to 0 uniformly in $t$, therefore we also have
$\lim_{q\to\infty}\mathbb{E}\int_{0}^{1}\log|{H}^{\pm}_{q}(t)|\mathds{1}_{\lvert{H}^{\pm}_{q}(t)\rvert\geq\varepsilon}\mathrm{d}t=\mathbb{E}\int_{0}^{1}\log|{G_{0.2,1.1}}(t)|\mathds{1}_{|{G_{0.2,1.1}}(t)|\geq\varepsilon}\mathrm{d}t.$
We conclude in exactly the same way as in the previous subsection, i.e. we have
(11) $\lim_{q\to\infty}\mathbb{E}\int_{0}^{1}\log\left|\frac{1}{1.1^{1/2}}{H}^{\pm}_{q}(t)\right|\mathrm{d}t=\mathbb{E}\int_{0}^{1}\log\left|\frac{1}{1.1^{1/2}}{G_{0.2,1.1}}(t)\right|\mathrm{d}t.$
To finish the proof of Corollary 2, note that the left-hand side of (11) is the limit of the normalized logarithmic Mahler measures of the Littlewood polynomials
$\sum_{n=1}^{\lfloor 1.1q\rfloor}\chi(n+\lfloor 0.2q\rfloor)x^{n-1}\pm x^{q-\lfloor 0.2q\rfloor-1}$
(upon noticing that Mahler measure is invariant under multiplication by powers of $x$).
@@ 9. Proof of Theorem 4
$F_{q,a}(t):=S\left(\chi_{q},\frac{a}{q},1,t\right)=\sum_{n\leq q}\left(\frac{n+a}{q}\right)e(nt).$
@@
$\lim_{q\to\infty}\frac{1}{\sqrt{q}}\|F_{q,a}\|_{2k}=\phi_{k}(\alpha)$
when $a/q\to\alpha$. We will show that $\phi_{k}(\alpha)$ achieves its minimum at $\alpha=1/4$ for every $k\in\mathbb{N}$. We start with the following lemma.
Lemma 13.
We have
(12) $\phi_{k}(\alpha)^{2k}=\frac{1}{(2\pi)^{2k}}\int_{0}^{1}\mathbb{E}\left\lvert(e(t)-1)\sum_{m\in\mathbb{Z}}\frac{e(m\alpha)}{m+t}\mathbb{Y}(m)\right\rvert^{2k}\mathrm{d}t,$
where the $\mathbb{Y}(m)$ are $\pm 1$-valued uniform random variables.
Proof.
Let $G_{q,\alpha}=G_{q,\alpha,1}$. By definition of $\phi_{k}$ and $G_{q,\alpha}$, we have
$\phi_{k}(\alpha)^{2k}=\lim_{q\to\infty}\mathbb{E}\int_{0}^{1}\left\lvert G_{q,\alpha}(t)\right\rvert^{2k}\mathrm{d}t.$
Moreover, for any large $M>0$,
$\mathbb{E}\int_{0}^{1}\left\lvert G_{q,\alpha}(t)\right\rvert^{2k}\mathrm{d}t=\mathbb{E}\min\left\{\int_{0}^{1}\left\lvert G_{q,\alpha}(t)\right\rvert^{2k}\mathrm{d}t,M\right\}\\ +O\left(\sum_{l=0}^{\infty}2^{l+1}M\mathbb{P}\left(2^{l}M\leq\int_{0}^{1}\left\lvert G_{q,\alpha}(t)\right\rvert^{2k}\mathrm{d}t\leq 2^{l+1}M\right)\right).$
The probability in the error term is of course bounded by
$\mathbb{P}\left(\int_{0}^{1}\left\lvert G_{q,\alpha}(t)\right\rvert^{2k}\mathrm{d}t\geq 2^{l}M\right).$
Markov’s inequality gives
$\mathbb{P}\left(\int_{0}^{1}\left\lvert G_{q,\alpha}(t)\right\rvert^{2k}\mathrm{d}t\geq 2^{l}M\right)\leq\frac{\mathbb{E}\left(\int_{0}^{1}\left\lvert G_{q,\alpha}(t)\right\rvert^{2k}\mathrm{d}t\right)^{2}}{2^{2l}M^{2}}\ll_{k}\frac{1}{2^{2l}M^{2}},$
since
$\mathbb{E}\left(\int_{0}^{1}\left\lvert G_{q,\alpha}(t)\right\rvert^{2k}\mathrm{d}t\right)^{2}\leq\mathbb{E}\int_{0}^{1}\lvert G_{q,\alpha}(t)\rvert^{4k}\mathrm{d}t\ll_{k}1.$
The last inequality follows, for instance, by the result of Günther and Schmidt that $\mathbb{E}\int_{0}^{1}\lvert G_{q,\alpha}(t)\rvert^{4k}\mathrm{d}t$ converges to $\phi_{2k}(\alpha)^{4k}$. We conclude that
$\mathbb{E}\int_{0}^{1}\left\lvert G_{q,\alpha}(t)\right\rvert^{2k}\mathrm{d}t=\mathbb{E}\min\left\{\int_{0}^{1}\left\lvert G_{q,\alpha}(t)\right\rvert^{2k}\mathrm{d}t,M\right\}+O\left(\frac{1}{M}\right).$
Now, $f\mapsto\min\left\{\int_{0}^{1}\left\lvert f(t)\right\rvert^{2k}\mathrm{d}t,M\right\}$ is a bounded continuous functional. Hence as $q\to\infty$, by Theorem 2 we have
$\phi_{k}(\alpha)^{2k}=\mathbb{E}\min\left\{\frac{1}{(2\pi)^{2k}}\int_{0}^{1}\left\lvert(e(t)-1)\sum_{m\in\mathbb{Z}}\frac{e(m\alpha)}{m+t}\mathbb{Y}(m)\right\rvert^{2k}\mathrm{d}t,M\right\}+O\left(\frac{1}{M}\right).$
Letting $M\to\infty$ and applying the monotone convergence theorem yields (12). ∎
Expanding the expectation in (12), the integral is equal to
$\int_{0}^{1}\lvert e(t)-1\rvert^{2k}\sum_{m_{i},n_{i}\in\mathbb{Z}}\Bigg(\frac{e(\alpha(m_{1}+\ldots+m_{k}-n_{1}-\ldots-n_{k}))}{(m_{1}+t)\cdots(m_{k}+t)(n_{1}+t)\cdots(n_{k}+t)}\times\\ \mathbb{E}\big[\mathbb{Y}(m_{1})\cdots\mathbb{Y}(m_{k})\mathbb{Y}(n_{1})\cdots\mathbb{Y}(n_{k})\big]\Bigg)\mathrm{d}t.$
@@
$\int_{0}^{1}\mathbb{E}\left\lvert(e(t)-1)\sum_{m\in\mathbb{Z}}\frac{e(m\alpha)}{m+t}\mathbb{Y}(m)\right\rvert^{2k}\mathrm{d}t=\\ \int_{0}^{1}\lvert e(t)-1\rvert^{2k}\sum_{r=0}^{k}\sum_{\begin{subarray}{c}m_{r+1},\ldots,m_{k}\in\mathbb{Z}\\ n_{2r+1},\ldots,n_{k}\in\mathbb{Z}\end{subarray}}\frac{\mathbb{E}\big[\mathbb{Y}(m_{2r+1})\cdots\mathbb{Y}(m_{k})\mathbb{Y}(n_{2r+1})\cdots\mathbb{Y}(n_{k})\big]}{(m_{2r+1}+t)\cdots(m_{k}+t)(n_{2r+1}+t)\cdots(n_{k}+t)}M_{2r,t}(\alpha)\mathrm{d}t,$
@@
Now writing $\theta=2\alpha$ and continuing from (13), we write the square of the absolute value as the sum of the squares of the real and imaginary parts to get
@@ Acknowledgements
We are grateful to Yu Chen Sun, Neo Tardy and Victor Wang for fruitful conversations. B.Sh. is funded by a University of Bristol PhD scholarship.